9. Find the general solution y(t) using the method of undetermined coefficients.(d) y 4y + 4y = tºe21
Answer #1

volution: +0 Given equation is y+ 4y+ 4y = x 22+ Homogenous equation of is y+ 4y+ 4y =0 -- is characteristic equation of=[20k++ (2 6+ 2A) + 24 +5]C.PL Now y(t) = 2 [20&+ +(264 2A)+ + 2C+57e44+ [+A++ 284 2A]e 25 = [4x4 + (4818A)t + 4C + 40+24]e